The intensity of a sound wave at a distance of 5 m from the point source is 0.8 Wm⁻² . What will the intensity be at a distance of 10 m from the source?


1
Expert's answer
2021-11-30T17:44:21-0500
"I=0.8\frac{5^2}{10^2}=0.2\frac{W}{m^2}"